OpenStudy (anonymous):

Write an equation of a line parallel to y = 2/3 x -5 in slope-intercept form that passes through the point (-6, -1).

OpenStudy (anonymous):

a. y= 2/3 x + 3 b. y= 2/3 x - 6 c. y= -3x - 1 d. y= 3x + 6

OpenStudy (anonymous):

a

OpenStudy (anonymous):

WHen two lines are parallel, their slopes are equal. In the equation y=mx+c , m is the slop. The given equation suggests that the slope of the required equation is 2/3 hence the required equation is y=2/3x+c, where c is to be determined. Putting x=-6 and y=-1 we get c=3 hence the required equation is y=2/3x+3

OpenStudy (anonymous):

@s2tasker2

OpenStudy (anonymous):

Okay thank you (:

OpenStudy (calculusfunctions):

The equation of a line in slope-intercept form is written as y = mx + b, where m is the slope and b is the y-intercept of the line.

OpenStudy (calculusfunctions):

If the line is parallel to the given line, then it has the same slope.

OpenStudy (calculusfunctions):

If it passes through a point (x1, y1), then substitute x = x1 and y = y1 and solve for b.